OpenStudy (anonymous):

okay so I need confirmation that there is something wrong with this question. This is a midterm review and I'm FREAKING out here. I'm not sure if i'm doing this all wrong, or what? Suppose y varies directly with x, and y=-14 when x=9. What direct variation equation relates x and y? A) Y=-.64x B)Y= 1.56x C)Y=-0.07x D) Y= -1.56x (the closest thing i got to -14 was answer D, as -14.04 ,PLEASE HELP ASAP!!!)

OpenStudy (anonymous):

No there is an answer, it's d. Check your work.

jimthompson5910 (jim_thompson5910):

"Suppose y varies directly with x" means that y = kx where k is some constant Since we are told that "y=-14 when x=9" we can plug these two pieces of info in and solve for k to get y = kx -14 = k*9 -14/9 = k k = -14/9 k = -1.56 (approximate and we're rounding to the nearest hundredth) So that explains why the answer is y = -1.56x

OpenStudy (anonymous):

y = kx -14 = 9k k = -14/9 k is roughly -1.55556 So it's: Y = -1.56x

OpenStudy (anonymous):

thanks guys! i was gonna die there i'm so glad i found this site